please help me solve this question the length of a rectangle is 7 more than the width. the perimeter is 58 what is the length and the width im not sure how to even solve it and what the equation would be
Answer by rfer(16322)

2x+2(x+7)=58
2x+2x+14=58
4x=44
x=11
x+7=18